OpenStudy (anonymous):

Algebra Question: Input in standard form the equation of the given line.The line that passes through (-1, -3) and (2, 1) Algebra Question: Input in standard form the equation of the given line.The line that passes through (-1, -3) and (2, 1) @Mathematics

OpenStudy (anonymous):

The slope is 4/3

OpenStudy (anonymous):

But it has to be on standard form

OpenStudy (anonymous):

y-b=m(x-a)

OpenStudy (anonymous):

Yeah and I got y=4/8-3

OpenStudy (anonymous):

3y=4x-5 is what i got.

OpenStudy (anonymous):

Or if you like y=4/3x-5/3

OpenStudy (anonymous):

It has to be in standard form which is Ax-By=C

OpenStudy (anonymous):

if you like, the 4x - 3y = 5

OpenStudy (anonymous):

Yeah I think that is correct... let me check.

OpenStudy (anonymous):

Yup, that is correct... do you have time for another

OpenStudy (anonymous):

Input in standard form the equation of the given line.The line with m = 1/2 and b = 3.

OpenStudy (anonymous):

Sure why not.

OpenStudy (anonymous):

I don't know how to do this.

OpenStudy (anonymous):

It looks easy though.

OpenStudy (anonymous):

Best I can get is x - 2y = a+6 but I assume theres a way to work out a I'm not seeing

OpenStudy (anonymous):

Might be better reposting that problem get some of the better guys on it.

OpenStudy (anonymous):

Alright thanks.